EVA Report #6 – February 7th

Crew 221 EVA Report 07 Feb 2020

EVA # 6

Author: Luis Monge (Engineer)

Purpose of EVA: Collection of microbial samples at Candor Chasma deeper into the canyon for geobiology experiments being conducted in the science dome. Collection of more regolith-type substrates for the scientific objectives of the ISRU project (growing plants in Martian soil).

Start time: 09:55 am

End time: 12:30 pm

Narrative: Today we planned our second long EVA. L.Monge, H.Kalucha and M.Grulich went out to drive towards Candor Chasma with the aim to collect samples mainly for microbiology projects. The Martian weather is currently very much in our favor and at some point during the EVA all participants complained about the heat. The sun shining high and there was almost no wind. M.Grulich took the lead on spirit to guide the team stopping on the way to check the maps and road conditions. Once near the destination, the EVA crew left the rovers.

Once at Candor Chasma, Crew Commander M.Grulich was looking for geology traces where microbes could be found. Samples at seven different locations were collected, and in one location it was dug to get a sample form at least 40cm deep into the ground.

While taking samples for microbiology, Crew Commander M.Grulich was very careful using extra nitrile gloves on top of the EVA gloves and sterile sample containers and little shovels that have had been cleaned prior to the EVA. On the way back to the rovers a bag of regolith was collected for Greenhab experiments. All objectives of the EVA have been accomplished.

Destination: Candor Chasma.

Coordinates (use UTM NAD27 CONUS): 519750/4251900 (Parking), Stream Bed connector; 520378/4251094

EVA Participants: Maria Grulich (Commander), Luis Monge (Engineer), Hemani Kalucha (Journalist)

Road(s) and routes per MDRS Map: North on Cow Dung Road (0110) to Galileo Road (1104), East along Galileo Road and park the rovers. Walk along Cactus Road towards Candor Chasma.

Mode of travel: Driving and walking

Vehicles used: Spirit and Curiosity

Commander Report – February 7th


Crew 221, MDRS

Crew 221 Commander Report 7FEB2020

Author: Maria Grulich, Commander

Title: First days

The crew 221 Earth2Mars settled in and started they science projects. The crew did not realize how much also the daily duties of cooking, cleaning and checking on the habitat takes time. Going on EVA starting to become a routine: checking the bags, the radios, comms-check, suiting up the crew members, depressurization and post-EVA and pre-EVA checklists.

Also, the crew cultural nights, sport sessions and meditation sessions coming along, even though we do not manage to do it every evening as it was planned beforehand. We make good progress on our science projects, which works better than expected and the crew’s cooking skills improve every day. The crew is very conservative with water usage, which results that the crew can take a 3min showers every 2-3 days which everyone is excited about. We are all caught up on the several surveys we have to fill and try to keep track of our samples and EVAs even though not all the time the GPS tracker works as expected. The Martian weather is very much in our favour and we hope that it will stay like this.

The first week the focus was put on microbiology and ISRU projects as they need time to develop. The following week will have a focus on the Human Robotic projects and technology demonstrations. Looking forward to more days on Mars.

Commander Grulich

MDRS Station

Operations Report – February 7th

Crew 221 Operations Report 07Feb2022

SOL: 5

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: off (at 8pm) waiting for SOC to reach 80% to be turned on

Hours run: 11:00

From what time last night: 8:00pm

To what time this morning: 7:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 84%

Diesel Reading – 98%

Station Propane Reading – 60%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 320 gals

Static to Loft Pump used – yes

Water in GreenHab – 232.5 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129.5

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 63%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 135.7

Beginning charge: 100%

Ending charge: 45%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? yes, for mail run by Atila

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational, lose cable on suit 7 fixed

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues watered with 4L. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators running, microbiology experiments going on and cultures are being grown in agar plates.

Summary of RAM operations: robot assembly is ongoing and expected to continue.

Summary of any observatory issues: nominal

Summary of health and safety issues: crew member had epistaxis, treated inmediately by crew doctor, HSO report submitted.

Questions, concerns and requests to Mission Support: every time the oven is turned on there is a heavy gas smell.

GreenHab Report – February 7th

Crew 221 GreenHab Report 07FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on.

Average temperatures: 23.75°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 10 gallons

Daily water usage for research and/or other purposes: 0.5 gallons

Water in Blue Tank: 232.5 gallons

Time(s) of watering for crops: 3

Changes to crops: Ginger has sprouted! The wilty tomatoes look much better today after a healthy watering. Butter lettuce also has bugs so I gave it a little spray with the pesticide. One of the strawberry plants has flowered!

Narrative: I performed a moisture check on all the plant beds to give me a better idea of their water needs. The lettuce plants take up water quite slowly, whereas the herbs and the tomatoes require at least two waterings a day. Luis and I also checked on our Blue Corn experiment. Unfortunately the Martian soil is basically like cement or clay when it gets wet and so the corn is likely struggling. We’ve halved the soil samples and mixed half with Earth soil to see if that makes a difference. Maria and I are also planning a mungbeans experiment, looking at different Martian soil to Earth ratios to see if we can grow anything in the next week. We will set that up tomorrow.

Harvest: NA

Support/supplies needed: NA

Other Notes: N/A


Science Report – February 7th

Crew 221 Science Report 07-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we continued our microbiology experiment. We observed further growth in bacteria in the dirty space suit sample. The bacteria was primitively identified as Bacillus subtilis. We also successfully made some Agar today. We added sugar, water, gelatin, and beef boullion to make the Agar. We then added the Agar to 12 petri dishes. All samples were weighed and put into the catalogue. Lastly, we did some preliminary pH soil testing on the samples.

2. ISRU project
The martian soil was cracked and inhospitable to plants, so we adjusted the ISRU experiment to be half martian soil and half earth soil, with all the same other parameters. Hopefully, the results will be kinder.

Sol Summary – February 7th

Crew 221 Sol Summary Report 07Feb-2020


Summary Title: It’s all about the Science

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol Summary: Crew 221 had another cultural night with Jessica Todd the Greenhab Officer, introducing the crew to Australian history, slang, the most dangerous and cutest animals in Australia and Vegemite, which not every crew member enjoyed. The main take-away is that in Australia everyone is your “mate” and you have to help them just like on Mars.

This morning, the commander woke up the team to prepare for another long EVA to go to Candor Chasma, which the crew expects to have been formed by water. The EVA team left at 10:00 am and had beautiful weather and the geology of Candor Chasma was very interesting. There were several stratigraphic layers, from which the crew took samples and will try to grow bacteria.

In the afternoon, the team worked on their robot under the eyes of “Fuzz” and made good progress besides one little piece broke off, which has to be glued before the team can continue and probably finish the robot tomorrow.

The samples in the incubator currently only show bacteria growth on the sample of the spacesuit before cleaning. This is also a good sign, since it means the crew is working cleanly. Today, they also produced agar, which was successful this time. All samples have been weighed and tomorrow, they will be processed to grow microbes. Tonight, there will be pizza prepared by the Crew Journalist H. Kalucha and we have a sports session with Crew Engineer L. D. Monge.

Look Ahead Plan: Sports Session, Meditation Session

Anomalies in work: None

Update mouse in hab: still no joy in catching it.

Weather: sunny no winds in the morning, windy in the afternoon, no precipitation

Crew Physical Status: Good

EVA: 10:00an to 12:50pm

Reports to be filed:

Sol Summary

Commander Report

Photo Report

EVA Report

EVA Request

Operations Report including HSO Report

GreenHab Report

Science Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site